OpenStudy (anonymous):

14.) An isosceles triangle has an area of 48 ft2. If the base is 12 ft, what is the length of each leg?

OpenStudy (anonymous):

I know the height is 8, but that wouldnt be the side too right?

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \rm 24=\frac{h \times 6}{2} }\) \(\LARGE\color{blue}{ \rm 48=h \times 6 }\) \(\LARGE\color{blue}{ \rm 8=h }\)

OpenStudy (solomonzelman):

Use the Pythagorean to find the missing legs.

OpenStudy (solomonzelman):

\[6^2+8^2=c^2~~~~~~~~~~~-->~~~~~~~~c=\sqrt{6^2+8^2}\]

OpenStudy (solomonzelman):

the little rest is on you.

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

Anytime !